5. Find the general solution of the non-homogenous differential equation. y – 7y +12y = ek
Answer #1

sol E. y 7y+12y=e oh Auxiliary egh is m2-7m +12 = 0 m-4m-3m +12=0 mm-4)-3m-4)=0 (m-4) (-3) =0 m = 3,4 44 .: C:F CF F cesh +2e